Suppose the call money rate is 4.5 percent, and you pay a spread of 2.5 percent over that. You buy 900 shares of stock at $80 per share. You put up $46,800. One year later, the stock is selling for $90 per share and you close out your position. What is your return assuming a dividend of $0.86 per share is paid?
Answer #1

Profit = (selling price+dividends-puchase price)*number of shares-(call money+spread)*(purchase price*number of shares-amount as margin)

=(90+0.86-80)*900-(0.045+0.025)*(80*900-46800)=8010

Return = profit/amount as margin

=8010/46800=17.12%